# CoPlay: Audio-agnostic Cognitive Scaling for Acoustic Sensing

### v1

@@
Abstract.
Acoustic sensing manifests great potential in various applications that encompass health monitoring, gesture interface and imaging by leveraging the speakers and microphones on smart devices. However, in ongoing research and development in acoustic sensing, one problem is often overlooked: the same speaker, when used concurrently for sensing and other traditional applications (like playing music), could cause interference in both making it impractical to use in the real world. The strong ultrasonic sensing signals mixed with music would overload the speaker’s mixer. To confront this issue of overloaded signals, current solutions are clipping or down-scaling, both of which affect the music playback quality and also sensing range and accuracy. To address this challenge, we propose CoPlay, a deep learning based optimization algorithm to cognitively adapt the sensing signal. It can 1) maximize the sensing signal magnitude within the available bandwidth left by the concurrent music to optimize sensing range and accuracy and 2) minimize any consequential frequency distortion that can affect music playback. In this work, we design a deep learning model and test it on common types of sensing signals (sine wave or Frequency Modulated Continuous Wave FMCW) as inputs with various agnostic concurrent music and speech. First, we evaluated the model performance to show the quality of the generated signals. Then we conducted field studies of downstream acoustic sensing tasks in the real world. A study with 12 users proved that respiration monitoring and gesture recognition using our adapted signal achieve similar accuracy as no-concurrent-music scenarios, while clipping or down-scaling manifests worse accuracy. A qualitative study also manifests that the music play quality is not degraded, unlike traditional clipping or down-scaling methods.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 30th Annual International Conference On Mobile Computing And Networking. ; 2024; Washington, D.C., USAccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methodsccs: Applied computing Sound and music computing
1. Introduction
@@
Figure 1. Our cognitively scaled sensing does not overload the mixer when there is concurrent music, better than simply clipping or down-scaling which leads to frequency distortion or weakens signal power.
@@
Acoustic sensing has been an active field of research in mobile computing community. Typically, an acoustic sensing system leverages the acoustic sensors on commercial-off-the-shell (COTS) smart devices. The system emits inaudible custom ultrasound signals from the speaker and then records the reflections using microphones transforming the device into an active SONAR system that is capable of capturing millimeter-level motion in the environment enabling various applications such as human gesture recognition and tracking (Wang et al., 2022; Koteshwara et al., 2023; Wang et al., 2016; Nandakumar et al., 2016), biosignal monitoring (Nandakumar et al., 2015; Wang et al., 2019; Song et al., 2020). Acoustic sensing offers unique advantages as a novel perception modality: 1) Unlike cameras, it works even in non-line-of-sight or poor light conditions, providing enhanced privacy preservation. 2) The speed of sound is many orders of magnitude slower than the speed of radio frequency (RF), allowing for millimeter-level sensing resolution with a significantly narrower bandwidth. 3) They harness the increasing number of speakers and microphones in IoT devices without necessitating hardware modifications. As such, acoustic sensing is promising in fields like Human-Computer Interaction (HCI) and health monitoring. The research community has been actively working on improving acoustic sensing systems in terms of accuracy, security, etc.
@@
However, one critical issue is often overlooked in ongoing research and developments in acoustic sensing: most systems assume the speaker solely plays the sensing signal. Yet, speakers are conventionally used for speech or music applications. Concurrently using audio applications when emitting sensing signals can cause an overload in the speaker’s mixer. To elaborate, a speaker mixer, or audio mixer, is a software component in the operating systems that mixes audio streams from different applications. To accomplish this, the mixer performs resampling, scaling, and addition to consolidate them into a single stream for output. If the value of the mixed signal exceeds the maximum numerical range supported by the DAC, we call this problem signal overload.
In consequence, the speaker mixer would take two possible actions: one is to clip off the overloaded part; another one is to down-scale the signal to ensure the maximum is within the allowable range, as shown in Fig. 1. Currently, these two strategies are widely adopted by operating systems. In Android, Windows, and Linux, they clip off the exceeding portion; IOS and MacOS scale down all the audio streams by a factor.  (Li et al., 2022a; Koteshwara et al., 2023) identified that both these methods will cause degraded quality of music or sensing. 1) Clipping results in distortion in the frequency domain. For example, if the user is playing music or answering a phone call, they will hear buzz sounds. The transmitted sensing signal is also distorted. Since the receiver-side ranging algorithms usually rely on accurate spectrum analysis, such as Fast Fourier Transform (FFT), the clipping will negatively affect the sensing performance. 2) Although down-scaling will not distort the frequency domain, it could significantly decrease the transmitted power of the signal. On the one hand, the volume of the music play goes low thus affecting its usability. On the other hand, down-scaled sensing signals are so weak that they can not propagate far enough in the environment. It is an important factor in the sensing performance as many sensing systems assume that they can use a large volume for the ultrasound to reach a longer detection range.
@@
Beyond identifying the essence of the concurrent music problem in  (Li et al., 2022a; Koteshwara et al., 2023), some work tangentially address this using closed-form solutions like echo cancellation (Koteshwara et al., 2023) or interleaving (Nandakumar et al., 2017). But they are limited to certain tasks or hurts the performance. In this work, we aim to propose an algorithmic solution to this problem that outperforms either closed-form methods in prior work or clipping and down-scaling in the current OS.
So, we propose CoPlay, a scaling algorithm, to intelligently adapt sensing signals in consideration of concurrent music applications and maintain optimal sensing and music playback. In specific, it can simultaneously
  • maximize the sensing magnitude within the available bandwidth left by the concurrent music to enable optimal sensing performance.
  • minimize any consequential frequency distortion.
Nevertheless, achieving these two goals simultaneously is hard because tweaking the signal amplitude in the time domain will cause noisy distortion in the frequency domain. For example, amplitude modulation(AM) will introduce noisy side bands beside the carrier frequency. While, in our case, dynamically scaling amplitude will cause even more frequency distortion and then compromise the sensing.
@@
Therefore, a deep learning algorithm is necessary to dynamically scale the signal, by formalizing these two objectives as an optimization problem. We name it cognitive scaling model, inspired by the concept of cognitive radar (Bell et al., 2015), where the transmission (Tx) signal is adaptively adjusted based on environmental changes. The perception-action cycles of cognition algorithms enable real-time adaptation of the radar. In our paper, we dynamically adjust the Tx ultrasound sensing signal according to the concurrent music. The suggested cognitive scaling algorithm runs in the speaker mixer and aims to diminish the risk of overload distortion while preserving the quality of both sensing and music.
@@
Specifically, the cognitive scaling neural network should work for agnostic music and various types of sensing signals. So, in this work, we examine our model with sine wave and frequency-modulated continuous wave (FMCW) chirp, with different types of concurrent audio, including piano, hot songs, bass, and conversational podcasts. To elaborate, sine and FMCW are the most common signals for acoustic sensing regardless of the receiver-side algorithms. After adapting these signals using CoPlay, the receiver-side algorithm should still work as before, as long as the transmitted signal is barely distorted and is known to the receiver-side algorithms, such as the Doppler effect, time of arrival (ToA), dechirping, etc. More specifically, the sine wave (Koteshwara et al., 2023; Wang et al., 2016; Mao et al., 2019) is simple, narrow-band, but susceptible to interference and noise and has limited range resolution. In contrast, acoustic sensing systems using FMCW chirps (Nandakumar et al., 2015; Zhou et al., 2018; Li et al., 2022b) are better correlated and have flexible range resolution. But it is more complex in frequency spectrum thus more challenging for our model to generate accurately. We had a workshop paper(cop, 2023)(anonymized) presenting preliminary results of only sine waves to embark on prior research conversation. But it couldn’t extend to FMCW signals. In contrast, CoPlay incorporates advanced algorithms with broader experiments, allowing it to process both sine and FMCW.
@@
To mitigate the above challenges, we design a cognitive scaling neural network with WaveNet backbone and custom layers. The model can take multiple types of sensing signal and agnostic concurrent music as input. The output is the cognitively scaled sensing signal. Specifically, we design a custom link function with music as bias, which acts as a filter to normalize the signal within dynamic bandwidth. Besides, we design a kernel to alleviate low-frequency noise stemming from the spectral bias and aliasing in convolution layers(Zhang, 2019; Rahaman et al., 2019). At last, we encode the above optimization goals in the loss terms to maximize the amplitude and minimize the frequency distortion.
To evaluate the model performance beyond numerical loss, we also conduct field studies to show how the adapted signal works in downstream tasks. The study of respiration detection and gesture recognition with 12 users shows that our adapted signals yield the same level of performance as the no-music scenarios. Besides, qualitative study manifests that our algorithms achieve better quality than clipped or down-scaled signals.
To summarize, our work has the following contributions.
@@
  • We formalize the problem of concurrent sensing and music as an optimization problem and build a deep-learning-based cognitive scaling algorithm for the mixer, which can enable optimal acoustic sensing performance without affecting the quality of concurrent agnostic music playback.
@@
  • We conduct an extensive evaluation of complex types of sensing signals including sine and FMCW and microbenchmark the model performance using various ablation studies and datasets.
@@
  • Beyond the analysis of model outputs, we deploy the algorithm in real-world field studies with 12 users. The quantitative and qualitative results show that our algorithm performs well in downstream tasks of respiration detection and gesture recognition, while clipping and downscaling hurt the performance.
In the rest of the paper, first, we detail the related work, problem definition, and model design. Then, we evaluate the model performance with ablation studies, along with quantitative and qualitative field studies. Finally, we address the limitations and future work. Finally, we introduce the related work of acoustic sensing and waveform modeling.
2. Related work
@@
Acknowledging this study as a pilot algorithmic solution to the concurrent music issue in acoustic sensing by framing it as an optimization problem, this section delves into a rundown of this field and associated methodologies. First, we overview the existing acoustic sensing research, noting some studies have offered preliminary investigations of the concurrent music issue, and some tangentially addressed it with closed-form fixes. Secondly, we elaborate on inspirational methods we drew from other fields, including non-autoregressive approaches in audio generation, and anti-aliasing techniques in neural networks for computer vision.
Acoustic sensing: Wireless perception using acoustics is a rising realm for applications such as health monitoring and human-computer interaction. It transforms mobile devices with speakers and microphones into a Sonar system. Typically, the system emits an inaudible ultrasound from the speaker; then the reflections from objects are captured by the microphone and subjected to analysis measuring the Doppler effect(Gupta et al., 2012; Koteshwara et al., 2023), time of flight (ToF) (Li et al., 2022b), direction of arrival (DOA) (Mao et al., 2019; Garg et al., 2021), etc. The signal is usually custom-modulated, such as single-tone sine waves or Frequency-modulated Continuous Wave (FMCW). For example, Amazon Echo (Koteshwara et al., 2023), LLAP (Wang et al., 2016), and  (Mao et al., 2019; Fu et al., 2022) uses high-frequency sine waves for presence detection, fingertips tracking, hand tracking and silence speech recognition. Cat (Mao et al., 2016), EchoPrint (Zhou et al., 2018), ApneaApp (Nandakumar et al., 2015) and Lasense  (Li et al., 2022b) transmit FMCW signal for localization, face or respiration detection. Besides, synthetic aperture radar (SAR) algorithms are widely used for acoustic imaging, as adapted in  (Mao et al., 2018; Wang et al., 2022).
@@
These acoustic sensing research usually focus on improving the sensing accuracy in the absence of concurrent music. However, this problem could cause distortion or lack of power in the received sensing signal. To elaborate, the distortion degrades the receiver-side sensing accuracy because the sensing depends on precise frequency; theoretically, it is a convolution property (when using Fast Fourier Transform (FFT) or correlation), covering methods of the phase-based/frequency-based, DOA and SAR. This underscores the reason behind our incorporation of FFT within the loss function. As for the external background noise, which is more often mentioned in these works, we can easily filter it out with a high-pass filter. However, internal noise—like music playing from the same speaker can bring about different problems, as highlighted in the introduction section.
@@
The concurrent music problem: Covertband (Nandakumar et al., 2017), a human activity detection system using acoustics, first proposed to interleave the concurrent music with the sensing signal; but this method hurts the sample rate exponentially, which degrades both music and sensing, especially ultrasound (17kHz+) that requires at least double sample rate (34kHz) to avoid aliasing. Both  (Li et al., 2022a; Koteshwara et al., 2023) underscored that the concurrent music distortion results from mixer overflow and clipping. The experience paper (Li et al., 2022a) validated the performance of their algorithm (Li et al., 2022b) under down-scaling, without proposing new solutions to the problem. But as we discussed above, scaling down causes additional issues. The Google patent  (Koteshwara et al., 2023) tried to remove the concurrent music distortions in the spectrum by acoustic echo cancellation (AEC) component, assuming motions are non-symmetric in the spectrum but distortions are. However, this method only applies to simple tasks of sensing such as Doppler-effect-based presence detection. Therefore, in our prior workshop paper(cop, 2023)(anonymized), we attempted to formalize this problem as optimization that is better than down-scaling or clipping. However, this preliminary paper was limited to sine waves but coundn’t extend to FMCW signals. In contrast, CoPlay introduces advanced techniques with broader and larger-scale experiments, enabling it to process both sine and FMCW. The increased complexity in frequency necessitates the utilization of CoPlay’s customized model.
Waveform modeling: Our objective is to address this concern in the mixer for agnostic concurrent audio or operating systems and various downstream sensing tasks. This involves 1) maintaining adequate detection range through a minimal power scaling-down and 2) minimal distortion in frequency. So, we formalize it as an optimization problem and let deep learning act as a window function. This problem is close to waveform modeling in the general audio domain.
Non-autoregressive (non-AR) raw audio generation is an active domain. Their sequence-to-sequence model architecture design for raw waveform has proved effective in many tasks such as including speech synthesis(Oord et al., 2016), sound source separation(Stoller et al., 2018), and music generation(Goel et al., 2022). Standard sequence modeling approaches like CNNs(Oord et al., 2016), RNNs(Mehri et al., 2016), and Transformers(Child et al., 2019) are adapted for waveform modeling. Particularly, WaveNet (Oord et al., 2016) and its variants(Engel et al., 2017; Peng et al., 2020) are fundamental backbone components in such architectures. In this work, we adopt this backbone and show the limitations and comparison of other architectures through experiments.
@@
Our task differs from common waveform modeling, which focuses on audible sound below a few hundred Hz. In contrast, ultrasonic sensing signals are 17kHz+, posing unique challenges due to neural networks’ spectral bias toward learning low-frequency features(Rahaman et al., 2019). Some studies attribute this bias to aliasing, caused by passing the activations through down-sampling layers like convolution and max pooling. Nevertheless, these investigations per spectral bias in neural networks are primarily in computer vision tasks(Zhang, 2019; Michaeli et al., 2023), where their definition of high frequency is much lower than the 17kHz+ ultrasonic waveform. Hence, we draw inspiration from them and customize our own kernels to alleviate this issue in ultrasound modeling particularly.
3. Method
As outlined in Fig. 1, our method is to dynamically scale the sensing signal rather than uniformly down-scaling or clipping. We build a cognitive scaling neural network that can 1) minimal frequency distortion and 2) minimal power down-scaling. This method provides optimal sensing performance and ensures concurrent agnostic audio applications remain unaffected. In this section, we start with interpreting the issue of dynamic scaling, then formalizing our objectives as an optimization problem, with a focus on sine wave and FMCW chirps as representative sensing signals. Following this, we introduce the model structure and its customized layers that we present as a cognitive scaling neural network.
3.1. Understanding the effect of dynamically scaling the signal

### v3

@@
Abstract
Acoustic sensing manifests great potential in various applications like health monitoring, gesture interface, by utilizing built-in speakers and microphones on smart devices. However, in ongoing research and development, one problem is often overlooked: the same speaker, when used concurrently for sensing and other traditional audio tasks (like playing music), could cause interference in both, making it impractical to use. The strong ultrasonic sensing signals mixed with music would overload the speaker’s mixer. To confront this issue of overloaded signals, current solutions are clipping or down-scaling, both of which affect the music playback quality, sensing range, and accuracy. To address this challenge, we propose CoPlay, a deep learning-based optimization algorithm to cognitively adapt the sensing signal and run in real-time. It can 1) maximize the sensing signal magnitude within the available bandwidth left by the concurrent music to optimize sensing range and accuracy and 2) minimize any consequential frequency distortion that can affect music playback. We design a custom model and test it on common types of sensing signals (sine wave or Frequency Modulated Continuous Wave FMCW) as inputs alongside various agnostic types of concurrent music and speech. First, we micro-benchmark the model performance to show the quality of the generated signals. Secondly, we conducted 2 field studies of downstream acoustic sensing tasks on 2 devices in the real world. A study with 12 users proved that respiration monitoring and gesture recognition using our adapted signal achieve similar accuracy as no-concurrent-music scenarios, whereas baseline methods of clipping or down-scaling manifest worse accuracy. A qualitative study also justifies that CoPlay leaves music untouched, unlike clipping or down-scaling that degrade music quality.
Index Terms:
@@ Abstract
I Introduction
@@
Figure 1: With concurrent music, our cognitive scaling avoids mixer overload, while clipping and down-scaling cause frequency distortion or weaken signal power.
Figure 2: Inordinate scaling in time domain causes distortion in frequency domain, taking Amplitude Modulation as an example.
@@
Acoustic sensing is an active research field, typically leveraging the built-in acoustic sensors on commercial-off-the-shelf (COTS) smart devices: it emits inaudible custom ultrasound signals from the speaker and then records the reflections using microphones. This transforms the device into an active SONAR system, capable of motion detection for various applications such as human gesture recognition and tracking [1, 2, 3, 4], biosignal monitoring [5, 6, 7]. Acoustic sensing offers unique advantages as a novel perception modality: 1) Unlike cameras, it works even in non-line-of-sight or poor light conditions, providing enhanced privacy preservation. 2) The speed of sound is many orders of magnitude slower than the speed of radio frequency (RF), allowing for millimeter-level sensing resolution with a significantly narrower bandwidth. 3) They harness the increasing number of built-in speakers and microphones in IoT devices without necessitating hardware modifications for sensing. It has been actively studied per improving accuracy, security, etc.
@@
However, one critical issue in ongoing research and developments is unaddressed: most systems assume the speaker solely plays the sensing signal. Yet, speakers are conventionally used for audio applications such as music playback. Concurrently running music and sensing can overload the speaker’s mixer. To elaborate, a speaker/audio mixer is a software component in the operating systems that mixes audio streams from different applications, via resampling, scaling, and adding them into a single output stream. If the mixed signal exceeds DAC’s maximum numerical range, overload occurs.
In consequence, the mixer would take two possible actions: clipping off the overloaded part or down-scaling the whole to stay within range, as in Fig. 2. Currently, these two strategies are widely adopted by operating systems: Android, Windows, and Linux uses clipping; IOS and MacOS use scaling.  [8, 2] identified that both methods will degrade the quality of music or sensing. 1) Clipping results in distortion in the frequency domain, leading to buzzing noise and affecting the performance of spectrum-based sensing algorithms such as FFT. 2) Down-scaling preserves frequency domain but significantly decreases transmitted signal power, lowering music play volume and weakening sensing detection range.
@@
Beyond identifying the essence of the concurrent music problem in  [8, 2], some work tangentially address this using closed-form solutions like echo cancellation [2] or interleaving [9]. But they are limited to certain tasks or hurts the signal quality and sensing performance. In this work, we aim to propose an algorithmic solution that outperforms prior closed-form methods. So, we propose CoPlay, a scaling algorithm, to dynamically adapt sensing signals in consideration of concurrent music applications and maintain optimal sensing and music playback. In specific, it can simultaneously 1) maximize magnitude within available bandwidth left by concurrent music, enabling optimal sensing performance 2) minimize any consequential frequency distortion.
@@
Achieving these both goals simultaneously is hard because tweaking the signal amplitude in time domain introduces frequency domain distortion. For example, amplitude modulation(AM) creates noisy side bands beside the carrier frequency. Our dynamic scaling of amplitude is worse. Therefore, a deep learning algorithm is necessary to dynamically scale the signal, by formalizing these two objectives as an optimization problem. We name it cognitive scaling model, inspired by cognitive radar [10], where the transmission signal(TX) is adaptively adjusted based on environmental changes via real-time perception-action cycles. In our paper, we dynamically adjust the Tx ultrasound according to the concurrent music, running in the speaker mixer to avoid overload while preserving both sensing and music quality.
@@
Specifically, the cognitive scaling neural network should handle agnostic music and various types of sensing signals (regardless of the receiver-side sensing algorithms). So, in this work, we examine our model with two common sensing signals, sine wave and FMCW chirp, along with four types of concurrent audio like piano and podcast. Sine wave is simple, narrow-band, but prone to interference, while FMCW chirps [5, 11, 12] offers better correlation and flexible range resolution but are more complex in frequency spectrum thus more challenging for our model to generate accurately. We had a workshop paper[13](anonymized) presenting preliminary results of only sine waves to embark on prior research conversation. But it couldn’t extend to FMCW signals. In contrast, CoPlay incorporates advanced algorithms with broader experiments, allowing it to process both sine and FMCW.
@@
To mitigate these challenges, we design a cognitive scaling neural network with custom layers, taking music and sensing as input and cognitively scaled sensing signal as output. Specifically, a custom link function with music as bias can filter and normalize inputs within dynamic bandwidth. A custom kernel further alleviates low-frequency noise stemming from the spectral bias and aliasing in convolution layers[14, 15]. The loss terms also integrate the two optimization goals to maximize the amplitude and minimize frequency distortion. To evaluate the model performance beyond numerical loss, we also conduct field studies in the real world to show how the adapted signal works in downstream tasks. The studies of respiration detection and gesture recognition with 12 users show that our adapted signals perform as well as in no-music scenarios. Besides, qualitative study manifests that we have better quality than clipped or down-scaled signals. To summarize, our work has the following contributions.
@@
  • We formalize the concurrent music issue as an optimization problem and design a cognitive scaling deep learning model for the mixer, ensuring optimal sensing performance and leaving concurrent agnostic audio untouched.
@@
  • Our extensive evaluation covers complex types of sensing signals including sine and FMCW and microbenchmarking under various ablation studies and datasets.
@@
  • Beyond model evaluation, we deploy the algorithm in real-world field studies with 12 users. The quantitative and qualitative results in downstream tasks of respiration detection and gesture manifest our great performance, while clipping and down-scaling hurt the performance.
II Related work
@@
As a pilot algorithmic solution to the concurrent music issue in acoustic sensing based on optimization, this section contains a rundown of this field and associated methodologies. First, we overview existing acoustic sensing research, noting some studies have offered preliminary investigations of the concurrent music issue, and some tangentially addressed it with closed-form fixes. Secondly, we elaborate on inspirational methods drew from other fields, including non-autoregressive approaches in audio generation, and anti-aliasing techniques in neural networks for computer vision.
@@
Acoustic sensing is typically subjected to analysis measuring the Doppler effect[16, 2], time of flight (ToF) [12], direction of arrival (DOA) [17, 18], etc. The signal is usually custom-modulated, such as sine waves or FMCW. For example, Amazon Echo [2], LLAP [3], and  [17, 19] uses high-frequency sine waves for presence detection, fingertips tracking, hand tracking and silence speech recognition. Cat [20], EchoPrint [11], ApneaApp [5] and Lasense  [12] transmit FMCW signal for localization, face or respiration detection. Besides, synthetic aperture radar (SAR) algorithms are widely used for acoustic imaging, as adapted in  [21, 1]. These research usually focus on improving the sensing accuracy without considering concurrent music. However, this problem could distort or weaken the received sensing signal, which hurts the phase-based/frequency-based sensing methods like DOA and SAR, specifically their convolution property (when using FFT or correlation). This underscores the reason behind our incorporation of FFT within the loss function. As for the external background noise, which is more often mentioned in these works, we can easily filter it out with a high-pass filter. However, internal noise—like music playing from the same speaker can bring unique challenges, as highlighted above.
@@
The concurrent music problem: Covertband [9], a human activity detection system using acoustics, first proposed to interleave the concurrent music with the sensing signal; but this method hurts the sample rate exponentially, which degrades both music and sensing, especially ultrasound (17kHz+) that requires at least double sample rate (34kHz) to avoid aliasing. Both  [8, 2] underscored that the concurrent music distortion results from mixer overflow and clipping. The experience paper [8] validated the performance of their algorithm [12] under down-scaling, without proposing new solutions to the problem. But as we discussed above, scaling down causes additional issues. The Google patent  [2] tried to remove the concurrent music distortions in the spectrum by acoustic echo cancellation (AEC) component, assuming motions are non-symmetric in the spectrum but distortions are. However, it only applies to simple tasks of sensing such as Doppler-effect-based presence detection. Therefore, our prior workshop paper[13] formalized this problem as optimization, which is better than down-scaling or clipping. However, this preliminary paper was limited to sine waves, excluding FMCW signals. In contrast, CoPlay introduces advanced techniques with broader experiments, enabling processing for both sine and FMCW. The increased complexity in frequency necessitates the utilization of CoPlay’s customized model.
@@
Waveform modeling: Our objectives involve simultaneously maximizing the magnitude and minimizing consequential frequency distortion. This needs to be formalized as an optimization problem, making deep learning a window function. So, it is close to waveform modeling in the general audio domain. Non-autoregressive (non-AR) raw audio generation is an active domain. Their sequence-to-sequence model architecture design for raw waveform has proved effective in many tasks such as including speech synthesis[22], sound source separation[23], and music generation[24]. Standard sequence modeling approaches like CNNs[22], RNNs[25], and Transformers[26] are adapted for waveform modeling. Particularly, WaveNet [22] and its variants[27, 28] are fundamental backbone components in such architectures. In this work, we adopt this backbone and show the limitations and comparison of other architectures through experiments. Moreover, our task differs from common waveform modeling: the latter focuses on audible sound below a few hundred Hz; while our ultrasonic sensing signals are 17kHz+, posing unique challenges due to neural networks’ spectral bias toward learning low-frequency features[15]. Some studies attribute this bias to aliasing, caused by passing the activations through down-sampling layers like convolution and max pooling. Nevertheless, these investigations per spectral bias in neural networks are primarily in computer vision tasks[14, 29], whose definition of high frequency is much lower than 17kHz. Inspired by these works, we design custom kernels to alleviate this issue in ultrasound modeling particularly.
III Method
In this section, we first interpret CoPlay’s challenges of dynamic scaling and formalize our objectives as an optimization problem. Then, we introduce the model with custom layers.
III-A Understanding the effect of dynamically scaling the signal